Juan goes to work averaging 32 mph because traffic conditions. He return home at 38mph. His total time driving is 2.3 hrs. Find how long it took for him to get from work to home
Answer by ankor@dixie-net.com(22740)

Juan goes to work averaging 32 mph because traffic conditions.
He return home at 38mph. His total time driving is 2.3 hrs.
Find how long it took for him to get from work to home
;
Let t = his time from work to home
then
(2.3-t) = his time from home to work
:
The dist there and back are the same.
Write a dist equation; dist = speed * time
:
time back = time there
38t = 32(2.3-t)
38t = 73.6 - 32t
38t + 32t = 73.6
70t = 73.6
t = 73.6/70
t = 1.05 hr or 1 + .05(60) = 1 hr 3 min
:
:
Check this; find the time there: 2.3 - 1.05 = 1.25 hrs
Find the dist each way, should be equal
1.25 * 32 = 40 mi
1.05 * 38 = 40 mi